# Bringing NeRFs to the Latent Space: Inverse Graphics Autoencoder

### v1

@@ 1 Introduction
Latent image representations are increasingly used in computer vision tasks. Most recent methods utilize latent representations for various tasks such as image generation (Rombach et al., 2022; Ramesh et al., 2021; Esser et al., 2021), and image segmentation (Long et al., 2015; Ohgushi et al., 2020), as they allow to represent images in a compact form which reduces their representation complexity. Nevertheless, their adoption for 3D tasks, such as Novel View Synthesis (NVS), remains limited due to the inherent incompatibilities between image latent spaces and 3D modeling methods. These incompatibilities stem from the lack of an underlying 3D geometry in latent spaces on the one hand, and the necessity of 3D-consistency among images for solving the inverse graphics problem on the other hand. In practice, encoding two 3D-consistent views of an object does not lead to 3D-consistent latent representations; instead, it produces latent images with features that vary inconsistently across views. This prohibits the direct application of scene learning methods such as Neural Radiance Fields (Mildenhall et al., 2020, NeRF) on such representations.
Furthermore, recent methods in 3D reconstruction have attempted to enable certain applications of latent spaces for NeRFs but also encountered latent 3D-inconsistency issues induced by Autoencoders (AE). They sidestepped this issue by implementing custom, scene-dependent “adapters” (Khalid et al., 2023) or “refinement layers” (Park et al., 2024) to correct renderings of neural scene representations into standard latent image representations. However, as these adapters are scene-dependent, they only enable a limited subset of applications. Thus, a universal “3D-aware” image latent space, analogous to the 2D image space, remains an under-explored area of research with promising potential, as it would enable interoperability with latent-based 2D methods.
@@
In this paper, we start by proposing a technique that enables learning NeRF models in the latent space. We present a training pipeline comprising two stages: Latent Supervision supervises NeRFs with latent image representations, and RGB Alignment aligns the learned latent scene with the RGB space. Yet, its application in a standard latent space results in sub-optimal latent NeRFs, due to their aforementioned incompatibilities. To address this, we draw inspiration from the relationship between the image space and 3D space (Fig. 1a) and introduce 3D-aware latent spaces, which augment regular image latent spaces by incorporating an underlying 3D geometry. We present an Inverse Graphics Autoencoder (IG-AE) that embeds a 3D-aware latent space (Fig. 1b). To achieve this, we apply 3D regularization on the latent space of an autoencoder by jointly training it with a synthetic set of latent 3D scenes. During this process, we align the encoded views with the renderings of the latent 3D scenes, which enforces 3D-consistency. Additionally, we propose an autoencoder preservation process that simultaneously autoencodes real and synthetic data while performing our training, which allows us to conserve the reconstructive performance of the autoencoder.
@@ 2 Related Work
While NeRFs (Mildenhall et al., 2020) were originally conceived to work in the image space, they have also been extended to work in other feature spaces (Tschernezki et al., 2022; Kobayashi et al., 2022). Notably, Latent NeRFs are extensions of NeRFs to the latent spaces of image autoencoders. They have been particularly explored in works targeting applications such as scene editing (Park et al., 2024; Khalid et al., 2023) and text-to-3D generation (Metzer et al., 2023), which are not directly feasible with regular NeRFs. To circumvent the incompatibilities between NeRFs and latent spaces, these works have resorted to special adapter layers that correct NeRF renderings into standard latent image representations. For novel view synthesis, Aumentado-Armstrong et al. (2023) employ hybrid NeRFs that are trained to simultaneously render both RGB and latent components. This is done to supervise the NeRF geometry during training, while only keeping latent components at inference, which enables good NVS performances and rendering speed-ups. In this work, we aim to train NeRFs that operate fully within the latent space. We address the incompatibility between NeRFs and latent spaces by tackling its root cause: the need for a inverse graphics autoencoders that preserves 3D consistency, or in other words, the need for 3D-aware latent spaces.
@@
Subsequent to the introduction of NeRF (Mildenhall et al., 2020), a multitude of NeRF models emerged to improve upon the original architecture by accelerating training times (Müller et al., 2022; Kerbl et al., 2023), improving rendering quality (Barron et al., 2021; Chen et al., 2022), as well as targeting other limitations (Fridovich-Keil et al., 2023; Martin-Brualla et al., 2021; Yu et al., 2021). With the escalation of NeRF methods, Nerfstudio (Tancik et al., 2023) emerged as a unified PyTorch (Paszke et al., 2019) framework in which NeRF models are implemented using standardized implementations, making it straightforward for researchers and practitioners to integrate various NeRF models into their projects. Today, most well-known NeRF techniques are implemented in Nerfstudio (Müller et al., 2022; Barron et al., 2021; Mildenhall et al., 2020; Kerbl et al., 2023), thus allowing for a friction-free experience when testing and comparing novel techniques. While Nerfstudio currently provides a streamlined approach for NeRFs, it limits the representation of 3D scenes to the natural color space, hence limiting both the research and development of Latent NeRFs. In this work, we propose an open-source extension of the Nerfstudio framework that supports training any Nerfstudio architecture in a custom latent space, thus facilitating our current as well as future research in this area. Furthermore, we integrate our IG-AE into this extension, which enables us to evaluate various current NeRF architectures on the task of latent scene learning.
@@ 3 Latent NeRF
Figure 2: Comparison of IG-AE and a standard AE. Encoding 3D-consistent images using an AE leads to 3D-inconsistent latent images. When trained on such latents, NeRF renderings present artifacts when decoded. IG-AE presents a 3D-aware latent space with 3D-consistent latent images. Latent NeRFs trained with our IG-AE omit these artifacts and more closely match the ground truth.
@@ 3.2 Training
Overall, we divide latent NeRF training into Latent Supervision and RGB Alignment. Latent Supervision is an accelerated approach that learns 3D structures in the latent space thanks to reduced image resolutions, and RGB alignment enhances NVS quality in the RGB space by using RGB supervision and a decoder fine-tuning, while still rendering the NeRF in the latent space. Fig. 2 (AE columns) illustrates the results of our latent NeRF training method when applied in a standard latent space. Latent NeRFs trained in a standard latent space with our method learn a coarse geometry. However, due to the aforementioned incompatibilities between latent spaces and NeRFs, the decoded renderings present artifacts in the RGB space. To address this, we shift focus to resolving these incompatibilities in the next section, and present an inverse graphics autoencoder embedding a 3D-aware latent that is compatible with learning latent NeRFs.
@@ 4.3 Autoencoder Preservation
To avoid overfitting on synthetic data, we additionally inject real images from Imagenet (Deng et al., 2009) into our training pipeline, on which we also define a reconstructive objective. We denote =$\mathcal{I}=\{I_{1},...,I_{L}\}$ the dataset of real images. We define the reconstructive loss on an image $I_{j}\in\mathcal{I}$ as follows:
@@
where $\lambda_{\mathrm{p}}$ and $\lambda_{\mathrm{TV}}$ are hyper-parameters, $\hat{I}_{j}=D_{\psi}(k_{j})$ is the reconstruction of $I_{j}$, and $k_{j}=E_{\phi}(I_{j})$ is the encoding of $I_{j}$. $\mathcal{L}_{\mathrm{p}}$ is a perceptual loss that refines the autoencoder reconstructions by aligning the reconstructed and original images in the feature space of a pre-trained network. $\mathcal{L}_{\mathrm{TV}}$ is a total variation loss that acts as a regularization term to encourage spatial smoothness by penalizing high-frequency variations in the latent images. We found that it leads to latent images that more closely reassemble our 3D-consistent latents. More details about TV are present in Appendix A. This is done to prevent the IG-AE from converging towards a dual-mode solution, where it is only 3D-consistent for synthetic scenes while functioning as a normal AE for real scenes.
@@ 4.4 IG-AE Training Objective
Overall, we present a joint training objective for our IG-AE that regularizes its latent space with 3D geometry while preserving its autoencoding performances. Our IG-AE training objective minimizes the following loss:
@@ 5 Experiments
We evaluate our method by first training an IG-AE on synthetic scenes from Objaverse, as described in Section 4. Subsequently, we train Nerfstudio models in our 3D-aware latent space on scenes from out-of-distribution datasets. We also train Nerfstudio models on held-out scenes from Objaverse (results in Section B.3). We compare the novel view synthesis performance of the Nerfstudio models when trained in our 3D-aware latent space (IG-AE), a standard latent space (AE), and the RGB space. Moreover, we evaluate the auto-encoding performances of our IG-AE on held-out real images. Finally, we assess our design choices by presenting an ablation study of our method, where we omit either our 3D-regularization components or our AE preservation components.
5.1 Nerfstudio Extension
We extend Nerfstudio to support latent scene learning in a code which we include in the supplementary material and will publicly release as open-source. Our extension omits the constraint of training models in the RGB space and allows to train Nerfstudio models in any feature space. Concurrently, we incorporate our latent NeRF training pipeline as well as IG-AE into Nerfstudio while adhering to its coding conventions, thereby enabling the training of any Nerfstudio model in our 3D-aware latent space. We modify the ray batching process to make it correspond to full images, where the default behavior uses randomly sampled pixels. This is necessary to be able to decode the rendered images without breaking the gradient flow. Beyond reproducibility purposes for this paper, we hope the proposed generic extension will facilitate research on latent NeRFs and their applications.
@@ 5.2 Datasets
For 3D-regularization, we adopt Objaverse (Deitke et al., 2023), a large dataset of diverse synthetic 3D objects. Objaverse provides meshes of synthetic objects, and thus presents no artifact in its scenes and no approximation errors in camera parameters. We utilize $N=500$ objects from Objaverse. Each object is rendered from $V=300$ views at a $128\times 128$ resolution. Each view is rendered from a randomly sampled camera pose on the sphere surrounding the object.
@@
For NeRF evaluations, we train NeRFs on held-out scenes from Objaverse, and on scenes from three out-of-distribution datasets: Shapenet Hats, Bags, and Vases (Chang et al., 2015). All scenes are rendered with the same resolution and number of views as Objaverse.
Table 1: Main Results on ShapeNet datasets. All results are obtained by training NeRF models in Nerfstudio using our Latent NeRF Training Pipeline, and are averaged over $4$ scenes from each dataset. Our 3D-aware latent space generalizes to out-of-distribution datasets and is more suited for latent NeRF training.
5.3 Training Details
IG-AE training details.
We adopt the pre-trained “Ostris KL-f8-d16” VAE (Burkett, 2024) from Hugging Face, which has a downscale factor $l=8$, and $c=16$ feature channels in the latent space. We apply a Total Variation (TV) regularization on our Tri-Planes with a factor $\lambda_{\mathrm{TV}}^{\mathrm{3D}}$ which prevents overfitting on train views (we exclude it from Eq. 8 for clarity). Analogously, we also apply TV regularization on encoded real images $k_{j}$. A detailed presentation of TV and its benefits for NVS can be found in Appendix A. Training IG-AE takes $60$ hours on 4×4$\times$4 × NVIDIA L4 GPUs. Our detailed hyper-parameter settings can be found in Section C.1.
Nerfstudio models training details.
We test the following models from Nerfstudio: Vanilla-NeRF (Mildenhall et al., 2020), Instant-NGP (Müller et al., 2022), TensoRF (Chen et al., 2022), and K-Planes (Fridovich-Keil et al., 2023). For each method, the Nerfstudio framework provides a proprietary loss $F_{\theta}$, as well as a custom training procedure that includes specific optimizers, schedulers and other method-specific components. To train a latent NeRF in Nerfstudio, we first train the chosen model for 10 0001000010\,00010 000 iterations to minimize $\mathfrak{L}_{\mathrm{LS}}$ using the method-specific optimization process. Subsequently, we continue the training with 15 0001500015\,00015 000 iterations of RGB alignment by minimizing $\mathfrak{L}_{\mathrm{align}}$. To account for the change of image representations, we modulate the learning rate of each method by a factor of $\xi_{\mathrm{LS}}$ in latent supervision, and a factor $\xi_{\mathrm{align}}$ for RGB alignment. Section C.2 details the hyper-parameters we used in Nerfstudio, including the values of these factors for each method.
@@ 5.4 Results
We report for each experiment the Peak Signal-to-Noise Ratio (PSNR $\uparrow$) for pixel-level similarity, the Structural Similarity Index Measure (SSIM $\uparrow$) for structural-level similarity, and the Learned Perceptual Image Patch Similarity (Zhang et al., 2018, LPIPS $\downarrow$) for perceptual similarity. Quantitative and qualitative results can be found in Tables 1 and 5, where we utilize our latent NeRF training pipeline to train various Nerfstudio models on out-of-distribution datasets from ShapeNet, both using our IG-AE and a standard AE. As illustrated, models trained with our IG-AE showcase a superior NVS performance compared to those trained with a standard AE. This proves that our IG-AE embeds a 3D-aware latent space that is more compatible with NeRF training than a standard latent space, and that is able to generalize to novel datasets.
Section B.1 presents the evolution of the NVS performance throughout our two Latent NeRF training stages. In Section B.2, we provide a comparison of latent NeRF training with classical RGB training. Latent NeRF training showcases significant speedups both in terms of training and rendering times, but lower NVS performance. This is due to a limitation in the representation of high frequencies when decoding latent NeRF renderings, as it can be seen in Fig. 5. We leave this to future work, as our approach is the first to propose a 3D-aware latent space and to assess its performance with various NeRF architectures.

### v2

@@ Abstract
1 Introduction
Latent image representations are increasingly used in computer vision tasks. Most recent methods utilize latent representations for various tasks such as image generation (Rombach et al., 2022; Esser et al., 2021), and image segmentation (Long et al., 2015; Ohgushi et al., 2020), as they allow to represent images in a compact form which reduces their representation complexity.
Similarly to 2D computer vision, leveraging latent image representations for 3D tasks would bring numerous advantages. First, training NeRFs on lower-resolution latent images would enable significant speed-ups in both rendering and training times. While efforts towards NeRF speed improvements have traditionally focused on working directly on NeRF architectures, this approach would instead tackle the space in which NeRF models are trained. This means that it would offer a broader impact across various NeRF models, extending beyond a single specific NeRF improvement. Second, latent NeRFs unlock an inter-operability between NeRFs and pre-trained latent-based methods. For instance, previous works have enabled this inter-operability in a task-specific, ad-hoc manner to unlock applications like scene editing (Park et al., 2024) and generation (Metzer et al., 2023).
@@ 1 Introduction
Despite the potential of latent NeRFs, the adoption of latent image representations for 3D tasks, such as Novel View Synthesis (NVS), remains limited due to the inherent incompatibilities between image latent spaces and 3D modeling methods. These incompatibilities stem from the lack of an underlying 3D geometry in latent spaces on the one hand, and the necessity of 3D-consistency among images for solving the inverse graphics problem on the other hand. In practice, encoding two 3D-consistent views of an object does not lead to 3D-consistent latent representations; instead, it produces latent images with features that vary inconsistently across views. This prohibits the direct application of scene learning methods such as NeRF (Mildenhall et al., 2020) on such representations.
To sidestep these incompatibilities, the aforementioned methods have resorted to the ad-hoc implementation of custom, scene-dependent “adapters” (Khalid et al., 2023) or “refinement layers” (Park et al., 2024). However, these solutions are only workarounds and do not address the root cause of the problem: the need for a universal “3D-aware” latent space induced by an image autoencoder that preserves 3D-consistency. Such an autoencoder would enable a broader compatibility between image autoencoders and NeRFs, unlocking the potential of latent NeRFs. In this paper, we explore this direction for the first time.
@@
We start by proposing a technique that enables learning NeRF models in the latent space. Our training pipeline comprises two stages: Latent Supervision supervises NeRFs with latent image representations, and RGB Alignment aligns the learned latent scene with the RGB space. Yet, its application in a standard latent space results in sub-optimal latent NeRFs, due to their aforementioned incompatibilities. To address this, we draw inspiration from the relationship between the image space and 3D space (Fig. 1a) and introduce 3D-aware latent spaces, which augment regular image latent spaces by incorporating an underlying 3D geometry. We present an Inverse Graphics Autoencoder (IG-AE) that embeds a 3D-aware latent space (Fig. 1b). To achieve this, we apply 3D regularization on the latent space of an autoencoder by jointly training it with a synthetic set of latent 3D scenes. During this process, we align the encoded views with the renderings of the latent 3D scenes, which enforces 3D-consistency. Additionally, we propose an autoencoder preservation process that simultaneously autoencodes real and synthetic data while performing our training, which allows us to conserve the reconstructive performance of the autoencoder.
@@ 2 Related Work
While NeRFs (Mildenhall et al., 2020) were originally conceived to work in the image space, they have also been extended to work in other feature spaces (Tschernezki et al., 2022; Kobayashi et al., 2022). Notably, Latent NeRFs are extensions of NeRFs to the latent spaces of image autoencoders. They have been particularly explored in works targeting applications such as scene editing (Park et al., 2024; Khalid et al., 2023) and text-to-3D generation (Metzer et al., 2023), which are not directly feasible with regular NeRFs. To circumvent the incompatibilities between NeRFs and latent spaces, these works have resorted to special adapter layers that correct NeRF renderings into standard latent image representations. For novel view synthesis, Aumentado-Armstrong et al. (2023) employ hybrid NeRFs that are trained to simultaneously render both RGB and latent components. This is done to supervise the NeRF geometry during training, while only keeping latent components at inference, which enables good NVS performances and rendering speed-ups. In this work, we aim to train NeRFs that operate fully within the latent space. We address the incompatibility between NeRFs and latent spaces by tackling its root cause: the need for an inverse graphics autoencoder that preserves 3D consistency, or in other words, the need for a 3D-aware latent space.
Scene embeddings.
While methods like Latent NeRFs and other feature fields modify the NeRF rendering space by replacing scene images with “feature images”, other approaches embed the entire scene information into a “scene embedding”. Recent works (Wang et al., 2023; Kosiorek et al., 2021; Sajjadi et al., 2022; Lan et al., 2024) focus on training encoders to transform scene information (e.g. images or text) into embeddings that fully encapsulate the 3D scene. NeRFs can then be obtained directly (without training) from such scene embeddings. In contrast, our approach aligns more closely with feature field methods (Tschernezki et al., 2022; Kobayashi et al., 2022), where NeRFs are trained to render “feature images” derived from scene images (e.g. DINO or CLIP features), enabling applications such as segmentation and object retrieval. However, we train NeRFs to render latent image representations obtained from an autoencoder fine-tuned for 3D tasks, to tackle the task of 3D reconstruction. For an elaborate distinction across recent works, we refer the reader to Appendix A.
@@
Subsequent to the introduction of NeRF (Mildenhall et al., 2020), a multitude of NeRF models emerged to improve upon the original architecture by accelerating training times (Müller et al., 2022; Kerbl et al., 2023), improving rendering quality (Barron et al., 2021; Chen et al., 2022), as well as targeting other limitations (Fridovich-Keil et al., 2023; Martin-Brualla et al., 2021; Yu et al., 2021). With the escalation of NeRF methods, Nerfstudio (Tancik et al., 2023) emerged as a unified PyTorch (Paszke et al., 2019) framework in which NeRF models are implemented using standardized implementations, making it straightforward for researchers and practitioners to integrate various NeRF models into their projects. Today, most well-known NeRF techniques are implemented in Nerfstudio (Müller et al., 2022; Barron et al., 2021; Mildenhall et al., 2020; Kerbl et al., 2023), thus allowing for a friction-free experience when testing and comparing novel techniques. While Nerfstudio currently provides a streamlined approach for NeRFs, it limits the representation of 3D scenes to the natural color space, hence limiting both the research and development of Latent NeRFs. In this work, we propose an open-source extension of the Nerfstudio framework that enables training supported NeRF methods in a custom latent space, thus facilitating our current as well as future research in this area. Furthermore, we integrate our IG-AE into this extension, which enables us to evaluate various current NeRF architectures on the task of latent scene learning.
@@ 3 Latent NeRF
Figure 2: Comparison of IG-AE and a standard AE. Encoding 3D-consistent images using an AE leads to 3D-inconsistent latent images. When trained on such latents, NeRF renderings present artifacts when decoded. IG-AE presents a 3D-aware latent space with 3D-consistent latent images. Latent NeRFs trained with IG-AE eliminate these artifacts and more closely match the ground truth.
@@ 3.2 Training
Overall, we divide latent NeRF training into Latent Supervision and RGB Alignment. Latent Supervision is an accelerated approach that learns 3D structures in the latent space thanks to reduced image resolutions, and RGB alignment enhances NVS quality in the RGB space by using RGB supervision and a decoder fine-tuning, while still rendering the NeRF in the latent space. Fig. 2 (AE columns) illustrates the results of our latent NeRF training method when applied in a standard latent space. Latent NeRFs trained in a standard latent space with our method learn a coarse geometry. However, due to the aforementioned incompatibilities between latent spaces and NeRFs, the decoded renderings present artifacts in the RGB space. To address this, we shift our focus to resolving these incompatibilities in the next section, and present an inverse graphics autoencoder embedding a 3D-aware latent that is compatible with learning latent NeRFs.
@@ 4.3 Autoencoder Preservation
To avoid overfitting on synthetic data, we additionally inject real images from Imagenet (Deng et al., 2009) into our training pipeline, on which we also define a reconstructive objective. We denote =$\mathcal{I}=\{I_{1},...,I_{L}\}$ the dataset of real images. We define the reconstructive loss on an image $I_{j}\in\mathcal{I}$ as:
@@
where $\lambda_{\mathrm{p}}$ and $\lambda_{\mathrm{TV}}$ are hyper-parameters, $\hat{I}_{j}=D_{\psi}(k_{j})$ is the reconstruction of $I_{j}$, and $k_{j}=E_{\phi}(I_{j})$ is the encoding of $I_{j}$. $\mathcal{L}_{\mathrm{p}}$ is a perceptual loss that refines the autoencoder reconstructions by aligning the reconstructed and original images in the feature space of a pre-trained network. $\mathcal{L}_{\mathrm{TV}}$ is a total variation loss that acts as a regularization term to encourage spatial smoothness by penalizing high-frequency variations in the latent images. We found that it leads to latent images that more closely reassemble our 3D-consistent latents. This is done to prevent the IG-AE from converging towards a dual-mode solution, where it is only 3D-consistent for synthetic scenes while functioning as a normal AE for real scenes. More details about TV are present in Appendix C.
@@ 4.4 IG-AE Training Objective
Overall, we present a joint objective for IG-AE that regularizes its latent space with 3D geometry while preserving autoencoding performances. The objective minimizes the following loss:
@@ 5 Experiments
We start by training an IG-AE on synthetic scenes from Objaverse, as described in Section 4. Subsequently, we train Nerfstudio models in our 3D-aware latent space on scenes from out-of-distribution datasets, as well as held-out scenes from Objaverse. To evaluate our method, we compare the novel view synthesis performance of the Nerfstudio models when trained in our 3D-aware latent space (IG-AE), a standard latent space (AE), and the RGB space. Moreover, we evaluate the auto-encoding performances of our IG-AE on held-out real images. Finally, we assess our design choices by presenting an ablation study of our method, where we omit either our 3D-regularization components or our AE preservation components. The details of our Nerfstudio extension, as well as IG-AE and latent NeRF trainings, are available in Appendix B.
Table 1: Main Results on ShapeNet datasets. All results are obtained by training NeRFs with our Latent NeRF Training Pipeline, and are averaged over $4$ scenes from each dataset. Our 3D-aware latent space generalizes to out-of-distribution datasets and is more suited for latent NeRF training.
@@ 5.1 Datasets
For 3D-regularization, we adopt Objaverse (Deitke et al., 2023), a synthetic dataset which is standard when large-scale and diverse 3D data is needed (Liu et al., 2023; Shi et al., 2024).
Objaverse provides meshes of synthetic objects, and thus presents no artifact in its scenes and no approximation errors in camera parameters. We utilize $N=500$ objects from Objaverse. Each object is rendered from $V=300$ views at a $128\times 128$ resolution.
@@
For NeRF evaluations, we utilize synthetic, object-level data as it aligns with the training domain. As such, we train NeRFs on held-out scenes from Objaverse, and on scenes from three out-of-distribution datasets: Shapenet Hats, Bags, and Vases (Chang et al., 2015). Note that, due to the challenges latent NeRFs face with simple object-level scenes, our evaluation focuses on such scenes, and excludes more complex or real scenes. We discuss this in Section 5.4.
@@ 5.2 Results
We report for each experiment the PSNR ($\uparrow$), the SSIM ($\uparrow$), and the LPIPS ($\downarrow$) (Zhang et al., 2018). Quantitative and qualitative results can be found in Tables 1 and 5, where we utilize our latent NeRF training pipeline to train various Nerfstudio models on out-of-distribution datasets from ShapeNet, both using our IG-AE and a standard AE. Models trained with IG-AE showcase superior NVS performance compared to those trained with a standard AE. This shows that IG-AE embeds a 3D-aware latent space better suited for NeRF training and capable of generalizing to new datasets. Additional results on held-out Objaverse scenes can be found in Section D.3. To further support this, Table 6 presents latent NVS metrics computed on latent images. Latent NeRFs showcase significantly better latent NVS performance with IG-AE, further confirming its 3D-awareness. Table 4 presents the NVS performance in the two stage of Latent NeRF training. In Table 8, we provide a comparison of latent NeRF training with classical RGB training. Latent NeRF training showcases speedups both in terms of training and rendering times, but lower NVS performance, which we discuss in Section 5.4.